# An extension of May’s Theorem to three alternatives: axiomatizing Minimax voting

### v2

@@ 6 Conclusion
By adding a compelling set of axioms to those of May’s Theorem, Theorem 3.8 uniquely determines how to vote on profiles with three alternatives (modulo tiebreaking): Minimax voting. Of course, we have not shown that these axioms are the normatively best axioms to augment May’s or that Minimax is the normatively best method of voting on three alternatives. However, we believe that Theorem 3.8 and axiomatic characterizations of other voting methods for three alternatives greatly facilitate the normative comparison of costs and benefits of methods of voting beyond two alternatives. Given recent proposals to implement top-three general elections using Minimax in the United States (Foley and Maskin 2024), such comparisons have not only theoretical but also political importance.
@@ Acknowledgements
For helpful comments, we thank Yifeng Ding, Mikayla Kelley, David McCune, Milan Mossé, Zoi Terzopoulou, Bill Zwicker, the two anonymous reviewers, and the audience at the AMMCS 2023 special session on Decisions and Fairness, where this work was first presented.
@@ Appendix A Minimax ties and refinements of Minimax
In the second case, Minimax MB agrees with Stable Voting (Holliday and Pacuit 2023b, Prop. 1.1). This case also shows that for three alternatives, Minimax MB satisfies only near immunity to spoilers, not immunity to spoilers (since $b$ wins without $a$ in the profile and beats $a$ head-to-head, immunity to spoilers requires $b$ to win in the full profile). [^17] Moreover, Minimax MB violates even near immunity to spoilers starting at four alternatives. Consider a Condorcet loser case with $n<k$, so Minimax MB selects only $c$; but now add another alternative $d$, whom $b$ and $c$ both beat head-to-head. If $b$’s margin over $c$ is sufficiently large, then although $c$ is the unique winner without $d$ and beats $d$ head-to-head, when $d$ joins, $b$ will be the unique Minimax MB winner, which violates near immunity to spoilers. To satisfy immunity to spoilers as well as the other axioms of Theorem 3.8 for three alternatives, one can consider the method that only properly refines Minimax in the ascending case when $n=m$, selecting $c$ in this case. On the other hand, a noteworthy feature of Minimax MB is the following.
@@
Minimax MB satisfies positive responsiveness as stated in Remark 2.2 for any number of alternatives.
@@
Suppose $a$ is among the Minimax MB winners in $\mathbf{P}$, so $a$ is among the Minimax winners with the greatest marginal Borda score. Further suppose $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter in $V(\mathbf{P})$ improving $a$’s position in their ballot, while nothing else changes. [^18] More precisely, there is some $b\in X(\mathbf{P})$ such that (i) $(a,b)\not\in\mathbf{P}(i)$ changes to $(a,b)\in\mathbf{P}^{\prime}(i)$ or (ii) $(b,a)\in\mathbf{P}(i)$ changes to $(b,a)\not\in\mathbf{P}^{\prime}(i)$, while for all alternatives $c\in X(\mathbf{P})\setminus\{a\}$, $(a,c)\in\mathbf{P}(i)$ implies $(a,c)\in\mathbf{P}^{\prime}(i)$, $(c,a)\not\in\mathbf{P}(i)$ implies $(c,a)\not\in\mathbf{P}^{\prime}(i)$, and for all $d\in X(\mathbf{P})\setminus\{a\}$, $(c,d)\in\mathbf{P}(i)$ if and only if $(c,d)\in\mathbf{P}^{\prime}(i)$. Then each margin of $a$ vs. another alternative is at least as great in $\mathbf{P}^{\prime}$ as in $\mathbf{P}$, while at least one such margin is strictly greater in $\mathbf{P}^{\prime}$; and no other alternative’s margin against another alternative increases. Thus, $a$ still has a minimal Minimax score in $\mathbf{P}^{\prime}$, and its marginal Borda score increases, while no other alternative’s marginal Borda score increases. It follows that $a$ is the unique Minimax MB winner in $\mathbf{P}^{\prime}$. ∎
@@
There is no voting method that satisfies anonymity, neutrality, positive responsiveness, and immunity to spoilers as in Remark 2.2 for all three-alternative profiles. However, Minimax MB does so for all three-alternative profiles with no zero margins between distinct alternatives.
@@
For the first claim, suppose for contradiction that there is such a method $F$. Consider the eight-voter profile $\mathbf{P}$ in which six voter submit each linear order of $\{a,b,c\}$, respectively, while the other two voters submit $cba$ and $bca$, respectively. Note that $\mathrm{Margin}\mathbf{P}(c,b)=0$ and $\mathrm{Margin}_{\mathbf{P}}(b,a)=\mathrm{Margin}_{\mathbf{P}}(c,a)=2$. By a standard argument using anonymity and neutrality, we have $F(\mathbf{P})=\{b,c\}$ and $F(\mathbf{P}_{-a})=\{b,c\}$. Now let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by a voter with the ranking $bac$ switching to $bca$. Then by positive responsiveness, $F(\mathbf{P}^{\prime})=\{c\}$. Since $\mathbf{P}^{\prime}_{-a}=\mathbf{P}_{-a}$, we have $b\in F(\mathbf{P}_{-a}^{\prime})$. Moreover, $\mathrm{Margin}_{\mathbf{P}^{\prime}}(b,a)=\mathrm{Margin}_{\mathbf{P}}(b,a)=2$, so $F(\mathbf{P}_{\mid\{a,b\}})=\{b\}$ by Theorem 2.3. Hence immunity to spoilers requires $b\in F(\mathbf{P}^{\prime})$, contradicting $F(\mathbf{P}^{\prime})=\{c\}$.

### v4

@@ 6 Conclusion
By adding a compelling set of axioms to those of May’s Theorem, Theorem 3.8 determines how to choose winners in profiles with three alternatives: use Minimax—or some refinement thereof when there are tied margins. Of course, we have not shown that these axioms are the normatively best axioms to augment May’s or that Minimax is the normatively best method of voting on three alternatives. However, we believe that Theorem 3.8 and axiomatic characterizations of other voting methods for three alternatives greatly facilitate the normative comparison of costs and benefits of methods of voting beyond two alternatives. Given recent proposals to implement top-three general elections using Minimax in the United States (Foley and Maskin 2024), such comparisons have not only theoretical but also political importance.
@@ Acknowledgements
For helpful comments, we thank Yifeng Ding, Mikayla Kelley, David McCune, Milan Mossé, Zoi Terzopoulou, Bill Zwicker, the two reviewers and editor for Social Choice and Welfare, and the audience at the AMMCS 2023 special session on Decisions and Fairness, where this work was first presented.
Compliance with Ethical Standards
The authors have no conflicts of interest.
@@ Appendix A Minimax ties and refinements of Minimax
In the second case, Minimax MB agrees with Stable Voting (Holliday and Pacuit 2023b, Prop. 1.1). This case also shows that for three alternatives, Minimax MB satisfies only near immunity to spoilers, not immunity to spoilers (since $b$ wins without $a$ in the profile and beats $a$ head-to-head, immunity to spoilers requires $b$ to win in the full profile). [^18] Moreover, Minimax MB violates even near immunity to spoilers starting at four alternatives. Consider a Condorcet loser case with $n<k$, so Minimax MB selects only $c$; but now add another alternative $d$, whom $b$ and $c$ both beat head-to-head. If $b$’s margin over $d$ is sufficiently large, then although $c$ is the unique winner without $d$ and beats $d$ head-to-head, when $d$ joins, $b$ will be the unique Minimax MB winner, which violates near immunity to spoilers. To satisfy immunity to spoilers as well as the other axioms of Theorem 3.8 for three alternatives, one can consider the method that only properly refines Minimax in the ascending case when $n=m$, selecting $c$ in this case. On the other hand, a noteworthy feature of Minimax MB is the following.
@@
Minimax MB satisfies positive responsiveness as stated in Remark 2.2 (Barberà 1977, Def. 7) for any number of alternatives.
@@
Suppose $a$ is among the Minimax MB winners in $\mathbf{P}$, so $a$ is among the Minimax winners with the greatest marginal Borda score. Further suppose $\mathbf{P}^{\prime}$ is obtained from $\mathbf{P}$ by one voter improving $a$’s position in their ballot, while nothing else changes. [^19] More precisely, there is some $b\in X(\mathbf{P})$ such that (i) $(a,b)\not\in\mathbf{P}(i)$ changes to $(a,b)\in\mathbf{P}^{\prime}(i)$ or (ii) $(b,a)\in\mathbf{P}(i)$ changes to $(b,a)\not\in\mathbf{P}^{\prime}(i)$, while for all alternatives $c\in X(\mathbf{P})\setminus\{a\}$, $(a,c)\in\mathbf{P}(i)$ implies $(a,c)\in\mathbf{P}^{\prime}(i)$, $(c,a)\not\in\mathbf{P}(i)$ implies $(c,a)\not\in\mathbf{P}^{\prime}(i)$, and for all $d\in X(\mathbf{P})\setminus\{a\}$, $(c,d)\in\mathbf{P}(i)$ if and only if $(c,d)\in\mathbf{P}^{\prime}(i)$. Then each margin of $a$ vs. another alternative is at least as great in $\mathbf{P}^{\prime}$ as in $\mathbf{P}$, while at least one such margin is strictly greater in $\mathbf{P}^{\prime}$; and no other alternative’s margin against another alternative increases. Thus, $a$ still has a minimal Minimax score in $\mathbf{P}^{\prime}$, and its marginal Borda score increases, while no other alternative’s marginal Borda score increases. It follows that $a$ is the unique Minimax MB winner in $\mathbf{P}^{\prime}$. ∎
@@
There is no voting method that satisfies anonymity, neutrality, positive responsiveness as in Remark 2.2, and immunity to spoilers for all three-alternative profiles. However, Minimax MB, which satisfies anonymity, neutrality, and positive responsiveness, satisfies immunity to spoilers for all three-alternative profiles with no zero margins between distinct alternatives.
@@
For the first claim, suppose for contradiction that there is such a method $F$. Consider the eight-voter profile $\mathbf{P}$ in which each of the six linear orders on $\{a,b,c\}$ is submitted by one voter, while the other two voters, say $i$ and $j$, submit $bca$ and $cba$, respectively. Note that $\mathrm{Margin}_{\mathbf{P}}(c,b)=0$ and $\mathrm{Margin}_{\mathbf{P}}(b,a)=\mathrm{Margin}_{\mathbf{P}}(c,a)=2$. We claim that $a\not\in F(\mathbf{P})$. Toward a contradiction, suppose $a\in F(\mathbf{P})$. Let $\mathbf{P}^{*}$ be obtained from $\mathbf{P}$ by voter $i$ switching from $bca$ to $bac$, while $j$ still submits $cba$. Then by positive responsiveness, $a\in F(\mathbf{P})$ implies $F(\mathbf{P}^{*})=\{a\}$. Let $\mathbf{P}^{**}$ be obtained from $\mathbf{P}^{*}$ by $i$ switching from $bac$ to $abc$ and $j$ switching from $cba$ to $cab$. Then by positive responsiveness, $F(\mathbf{P}^{*})=\{a\}$ implies $F(\mathbf{P}^{**})=\{a\}$. On the other hand, since $\mathbf{P}^{*}$ consists of each of the six linear orders plus an extra copy of each of $bac$ and $cba$, while $\mathbf{P}^{**}$ consists of each of the six linear orders plus an extra copy of each of $abc$ and $cab$, so that the roles of $a$ and $b$ are exchanged from $\mathbf{P}^{*}$ to $\mathbf{P}^{**}$, $F(\mathbf{P}^{*})=\{a\}$ implies $F(\mathbf{P}^{**})=\{b\}$ by a standard argument using anonymity and neutrality. Thus, we have a contradiction, so $a\not\in F(\mathbf{P})$. Then since $b$ and $c$ play symmetrical roles in $\mathbf{P}$, it follows by standard arguments using anonymity and neutrality again that $F(\mathbf{P})=\{b,c\}$ and $F(\mathbf{P}_{-a})=\{b,c\}$. Now let $\mathbf{P}^{\prime}$ be obtained from $\mathbf{P}$ by a voter with the ranking $bac$ switching to $bca$. Then by positive responsiveness, $F(\mathbf{P}^{\prime})=\{c\}$. Since $\mathbf{P}^{\prime}_{-a}=\mathbf{P}_{-a}$, we have $b\in F(\mathbf{P}_{-a}^{\prime})$. Moreover, $\mathrm{Margin}_{\mathbf{P}^{\prime}}(b,a)=\mathrm{Margin}_{\mathbf{P}}(b,a)=2$, so $F(\mathbf{P}_{\mid\{a,b\}})=\{b\}$ by Theorem 2.3. Then since $a\not\in F(\mathbf{P}^{\prime})$, immunity to spoilers requires $b\in F(\mathbf{P}^{\prime})$, contradicting $F(\mathbf{P}^{\prime})=\{c\}$.